# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502 Japan

### v2

@@ 5.2 Intersecting points of the Binder parameter
Now, we can calculate the intersection points $(\beta_{\times},U_{\times})$, using the interpolating function (30) for two different lattice sizes $L_{1}$ and $L_{2}$. First, we solve the quadratic equation, $\tilde{U}(\beta)|_{L_{1}}=\tilde{U}(\beta)|_{L_{2}}$, and obtain $\beta_{\times}$ as a solution. Then, substituting the value of $\beta_{\times}$ into $\tilde{U}(\beta)|_{L_{1}}$, we obtain the value of $U_{\times}$ as well. The results of the intersection points for $L_{1}<L_{2}$ close to the critical point are shown on the $\beta\,\text{-}\,U$ plane in Figure 4. The data points corresponding to the same $L_{1}$ are denoted by the same symbols as $L$ in Figure 3.
Figure 4: The intersecting positions $(\beta_{\times},U_{\times})$ of the Binder parameter for the two lattice sizes $L_{1}$ and $L_{2}$ with $L_{1}<L_{2}$ are plotted on the $\beta\,\text{-}\,U$ plane for $\lambda=0$ (left) and $8$ (right). Here only the regions near the critical points are shown. The data points with the same $L_{1}$ are denoted by the same symbols as $L$ shown in the legend of Figure 3. Among the data points with the same $L_{1}$, those with the larger $L_{2}$ appear further to the bottom left, approaching the critical point $(\beta_{c},U^{*})$.
@@
5.3 Determination of the critical temperature
@@
Now, we determine the critical temperature $\beta_{c}$ by using the scaling property of intersections of the Binder parameter $U$ since the explicit value of $\beta_{c}$ is necessary to compute the critical exponents. We perform the simultanuous fit of $(\beta_{\times},U_{\times})$ by using Eqs. (28) and (29) to find optimal parameters $(\beta_{c},U^{*},y_{t},\omega,c_{\nu},c_{\omega})$.
$\lambda$ $\beta_{c}$ $U^{*}$ $y_{t}$ $\omega$ $c_{\nu}$ $c_{\omega}$ $\chi^{2}/\mathrm{dof}$ $L_{\mathrm{min}}$
0 0.693035(11) 0.6210(2) 1.42(1) 1.4(2) 0.04(1) 0.043(1) 1.22 8
8 0.319844(18) 0.4066(69) 1.78(1) 0.9(2) 0.31(6) 0.076(2) 1.39 10
@@
Table 2: The fitting results of the intersection points $(\beta_{\times},U_{\times})$ for $\lambda=0$ and $8$.
@@
The fitting results and the choice of $L_{\mathrm{min}}$ are summarized in Table 2. Here, we introduce $L_{\mathrm{min}}$ to remove the small $L_{1}$ data which suffers from a sizable finite-size effect. Thus, we take $L_{\mathrm{min}}\leq L_{1}<L_{2}$ in our analysis. We choose $L_{\mathrm{min}}$ so that $\chi^{2}/\mathrm{dof}$ is close to $1$ and the fitting error is small. Furthermore, we put $L_{\mathrm{diff}}$ to be $L_{2}-L_{1}\geq L_{\mathrm{diff}}$. If $L_{\mathrm{diff}}$ is too small, the slopes of the functions $\tilde{U}(\beta)|_{L_{1}}$ and $\tilde{U}(\beta)|_{L_{2}}$ are almost the same. In this case, the intersection point cannot be determined precisely. Here, we fix $L_{\mathrm{diff}}=4$ in the fitting, which removes the closest pairs of $L_{1}$ and $L_{2}$.
@@
Figure 5: The fitting results of the scaling behavior of the intersections $(\beta_{\times},U_{\times})$ are shown for $\lambda=0$ (left) and $8$ (right). The numerical results of $\beta_{\times}$ (top) and $U_{\times}$ (bottom) are plotted against the functions $f(L_{1},L_{2})$ and $g(L_{1},L_{2})$, respectively, where the best-fit values of $y_{t}$ and $\omega$ are used to define these functions. Note that the data points are aligned on straight lines, which indicates the validity of the fitting ansatz. Here, the data symbols denote the same ones by replacing $L$ to $L_{1}$ shown in the legend of Figure 3.
@@
The data of $\beta_{\times}$ and $U_{\times}$ are shown in the left and right panels of Figure 5, respectively. In these panels, we take the horizontal axis as $f(L_{1},L_{2})$ and $g(L_{1},L_{2})$ defined in Eq. (27) so that the linearity should indicate our finite-size scaling ansatz is working well. Here, to compute $f(L_{1},L_{2})$ and $g(L_{1},L_{2})$, we use the best-fit values of $y_{t}$ and $\omega$ in Table 2. The numerical data seen in Figure 5 shows visible linearity for both $\beta_{\times}$ and $U_{\times}$ with $\lambda=0$ and $8$, suggesting the validity of our ansatz.
@@
Figure 6: Comparison of the critical temperature $\beta_{c}$ of the Heisenberg model obtained in this paper and the previous studies by the Monte Carlo Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz .
@@
Given the success of the finite-size scaling ansatz, we obtain the critical temperatures for the two cases of $\lambda$ from the fitting results as
$\beta_{c}=\begin{cases}0.693035(11)&(\lambda=0),\\ 0.319844(18)&(\lambda=8).\end{cases}$ (31)
We use these values of $\beta_{c}$ in the following analyses. Our result of $\beta_{c}$ for $\lambda=0$ is compared with the previous Monte Carlo studies on the Heisenberg model in Figure 6.
5.4 Critical exponent: $\nu$
@@
Figure 7: The gradient of the Binder parameter at the critical temperature $\beta_{c}$ is obtained using the interpolating function (36) as $\tilde{U}^{\prime}(\beta_{c})$ and plotted against $L$ in log-log scale for $\lambda=0$ (left) and $8$ (right).
@@
based on the scaling property in Eq. (24). Here we include the term of $L^{-\epsilon}$ as a finite-size correction, which should appear when $L$ is small, following a similar analysis in Ref. Campostrini:2000iw ; 2019PhRvB.100v4517H ; Hasenbusch:2020pwj . In principle, the correction exponent $\epsilon$ should be related to the other critical exponents such as $\omega$ (see Ref. 2019PhRvB.100v4517H ), but the precise determination of $\epsilon$ from the fits requires more precisions of data in larger $L$ asymptotics. In this paper, as did in Refs. Campostrini:2000iw ; 2019PhRvB.100v4517H ; Hasenbusch:2020pwj , we fix the value of $\epsilon$ beforehand and estimate the optimal value of $\epsilon$ by the stability of the resulting fit.
@@
Figure 8: The results of $\nu$ obtained via the fitting ansatz (33) are plotted against the lower bound $L_{\mathrm{min}}$ of the fitting range for $\lambda=0$ (left) and $8$ (right). In both panels, the symbols of plus-blue, cross-green, circle-red, and square-purple depict the results with $\epsilon=\infty$, $2.0$, $3.0$, and $6.0$, respectively. The horizontal lines depict the results of the constant fitting with the error bands.
@@
$\nu=0.712(1)\begin{pmatrix}+3\\ -0\end{pmatrix}\qquad\text{for }\lambda=0$ (34)
for the Heisenberg model.
@@
$\nu=0.601(2)\begin{pmatrix}+0\\ -2\end{pmatrix}\qquad\text{for }\lambda=8$ (35)
@@
These results of $\nu$ above are compared with the results of previous studies with various methods as shown in Figure 1. Here, we compare the values of $\nu$ for $\lambda=0$, $8$, and $\infty$ obtained by the Monte Carlo Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz , the conformal bootstrap Chester:2020iyt , the $\varepsilon$-expansion Aharony_1 ; 2022NuPhB.98515990K , and the functional renormalization group Nakayama:2023wrx . For reference, the conformal bootstrap predicts $\nu=0.7117(2)$ 444 Here, the error in the bootstrap bound is not one-sigma range but rigorous Chester:2020iyt . and the latest Monte Carlo study Hasenbusch:2020pwj reports $\nu=0.71164(10)$ for the Heisenberg model. Our result at $\lambda=0$ is consistent with them. On the other hand, the result at $\lambda=8$ shows a large deviation from the analytical prediction by the 1-loop $\varepsilon$-expansion. Here, we note that the difference between the 1-loop result of the $\varepsilon$-expansion and the 3-loop result at $\lambda=0$ is larger than the one between the 1-loop one at $\lambda=0$ and $\lambda=8$. It indicates that loop corrections are sizable, at least larger than the $\lambda$-dependence.
Furthermore, let us discuss the self-consistency of our results. As shown in Table 2, we obtained the estimations of $y_{t}=1/\nu$ by analysing the intersections of the Binder parameter in Section 5.3, which yields $\nu=0.706(5)$ and $0.563(4)$ for $\lambda=0$ and $8$, respectively. Compared with these values, the results of $\nu$ obtained in this section have smaller statistical errors. The smaller statistical error is mainly because we have fewer parameters to fit by considering the derivative of the Binder parameter and fixing $\beta_{c}$. The results of the two methods agree within the error for $\lambda=0$ but not well for $\lambda=8$. One of the reasons for this discrepancy could arise mainly from the choice of $L_{\mathrm{min}}$, where we set $L_{\mathrm{min}}=10$ in Section 5.3 and $L_{\mathrm{min}}\geq 16$ in this section.
5.5 Critical exponents: $\eta$ and $\gamma$
@@
Figure 10: The magnetic susceptibility at the critical temperature is obtained using the interpolating function (36) as $\tilde{\chi}_{m}(\beta_{c})|_{L}$ and plotted against $L$ in log-log scale for $\lambda=0$ (left) and $8$ (right).
@@
Figure 11: The results of $\eta$ obtained via the fitting ansatz (37) are plotted against the lower bound $L_{\mathrm{min}}$ of the fitting range for $\lambda=0$ (left) and $8$ (right). In both panels, the symbols of, plus-blue, cross-green, and circle-red depict the results with $\epsilon=\infty$, $1.0$, and $2.0$, respectively. The square-purple symbols correspond to $\epsilon=4.0$ in the left panel and $6.0$ in the right panel. The colored horizontal lines depict the results of the constant fitting with the error bands.
We compute the critical exponent $\eta$ based on the scaling property (25). Adding the finite size correction term, we consider the scaling function to be
@@
We take the result of $\epsilon=2.0$ as the central value and use the others to estimate the systematic errors, which results in
$\eta=0.0318(27)\begin{pmatrix}+10\\ -7\end{pmatrix}\qquad\text{for }\lambda=0$ (38)
for the Heisenberg model. For reference, the conformal bootstrap predicts $\eta=0.0379(1)$ Chester:2020iyt . Here, the error in the bootstrap bound is not one-sigma range but rigorous.
@@
$\eta=0.132(8)\begin{pmatrix}+0\\ -4\end{pmatrix}\qquad\text{for }\lambda=8$ (39)
@@
$\gamma=\begin{cases}1.4009(32)\begin{pmatrix}+58\\ -0\end{pmatrix}_{\nu}\begin{pmatrix}+5\\ -7\end{pmatrix}_{\eta}&(\lambda=0),\\ 1.123(6)\begin{pmatrix}+0\\ -3\end{pmatrix}_{\nu}\begin{pmatrix}+2\\ -0\end{pmatrix}_{\eta}&(\lambda=8).\end{cases}$ (40)
Here, the first parentheses represent the statistical error while the second and third ones denote the systematic errors from $\nu$ and $\eta$, respectively.
Figure 12: The critical exponent $\gamma$ obtained in this paper and the previous studies by the Monte Carlo Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz (colored open symbols), the conformal bootstrap Chester:2020iyt (star-yellow), the $\varepsilon$-expansion 2022NuPhB.98515990K (filled symbols) are compared. The filled circle-black symbols depict the 1-loop results of the $\varepsilon$-expansion in Ref. Aharony_1 and the filled square-circle ones are the 3-loop results in Ref. 2022NuPhB.98515990K .
Our results of $\gamma$ are compared with the results of the previous Monte Carlo studies Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz , the $\varepsilon$-expansion Aharony_1 ; 2022NuPhB.98515990K , and conformal bootstrap Chester:2020iyt in Figure 12. The 3-loop calculations yield coincidental results $\gamma=1.380(8)$ and $1.381(8)$ for the Heisenberg model ($\lambda=0$) and the Heisenberg-dipolar model with $\lambda=\infty$ Kudlis_2022 ; Gimenez-Grau:2023lpz . Also, the conformal bootstrap predicts $\gamma=1.3964(5)$ Chester:2020iyt . On the other hand, our numerical results of $\gamma$ show a clear difference between $\lambda=0$ and $8$. This observation suggests a possibility that the difference between the critical exponents at the dipolar and Heisenberg fixed points is more significant than the expectation from the 3-loop calculation. Although our numerical method has room to be improved for more rigorous investigation, our result is consistent with the previous studies at least for $\lambda=0$. For instance, the latest one, $\gamma=1.39635(20)$, obtained by the Monte Carlo study Hasenbusch:2020pwj is $2\sigma$ consistent (see the detailed comparison in Figure 12).
6 Correlation functions and transverse suppression

### v3

@@ V.2 Intersecting points of the Binder parameter
Now, we can calculate the intersection points $(\beta_{\times},U_{\times})$, using the interpolating function (30) for two different lattice sizes $L_{1}$ and $L_{2}$. First, we solve the quadratic equation, $\tilde{U}(\beta)|_{L_{1}}=\tilde{U}(\beta)|_{L_{2}}$, and obtain $\beta_{\times}$ as a solution. Then, substituting the value of $\beta_{\times}$ into $\tilde{U}(\beta)|_{L_{1}}$, we obtain the value of $U_{\times}$ as well. The results of the intersection points, $\beta_{\times}$ and $U_{\times}$, for $L_{1}<L_{2}$ are plotted against $1/L_{2}$ in Figure 4. The data points corresponding to the same $L_{1}$ are denoted by the same symbols as $L$ in Figure 3.
Figure 4: The intersecting positions $\beta_{\times}$ (top) and $U_{\times}$ (bottom) of the Binder parameter for the two lattice sizes $L_{1}$ and $L_{2}$ with $L_{1}<L_{2}$ are plotted against $1/L_{2}$ for $\lambda=0$ (left) and 88 (right). The data points with the same $L_{1}$ are denoted by the same symbols as $L$ shown in the legend of Figure 3. The data points converge to the critical values, $\beta_{c}$ and $U^{*}$, as both $L_{1}$ and $L_{2}$ increase.
@@
V.3 Determination of the critical temperature
@@
Now, we determine the critical temperature $\beta_{c}$ by using the scaling property of intersections of the Binder parameter $U$ since the explicit value of $\beta_{c}$ is necessary to compute the critical exponents. We perform the simultaneous fitting of $(\beta_{\times},U_{\times})$ by using Eqs. (28) and (29) to find optimal parameters $(\beta_{c},U^{*},y_{t},\omega,c_{\nu},c_{\omega})$.
$\lambda$ $L_{\mathrm{min}}$ $\beta_{c}$ $U^{*}$ $y_{t}$ $\omega$ $c_{\nu}$ $c_{\omega}$ $\chi^{2}/\mathrm{dof}$
6 0.693015(8) 0.6206(1) 1.42(1) 1.1(1) 0.020(2) 0.043(1) 1.51
0 8 0.693035(11) 0.6210(2) 1.42(1) 1.4(2) 0.035(1) 0.043(1) 1.22
10 0.693054(13) 0.6213(2) 1.39(1) 2.1(4) 0.128(101) 0.047(2) 1.03
8 0.319887(9) 0.417(2) 1.85(1) 1.2(1) 0.52(5) 0.061(2) 4.11
8 10 0.319844(18) 0.407(7) 1.78(1) 0.9(2) 0.31(6) 0.076(2) 1.39
12 0.319803(35) 0.378(39) 1.74(1) 0.5(3) 0.20(2) 0.084(3) 0.97
@@
Table 2: The fitting results of the intersection points $(\beta_{\times},U_{\times})$ with the different choices of $L_{\mathrm{min}}$ for $\lambda=0$ and $8$. In these results, $L_{\mathrm{diff}}=4$ and $L_{\mathrm{max}}=40$ are fixed.
@@
In this analysis, we introduce $L_{\mathrm{diff}}$, $L_{\mathrm{min}}$, and $L_{\mathrm{max}}$ to control the fitting range as $L_{2}-L_{1}\geq L_{\mathrm{diff}}$ and $L_{\mathrm{min}}\leq L_{1}<L_{2}\leq L_{\mathrm{max}}$. The fitting results are summarized in Table 2, where we consider three different choices of $L_{\mathrm{min}}$ for each $\lambda$. We fix $L_{\mathrm{diff}}=4$ to remove the closest pairs of $L_{1}$ and $L_{2}$ from the fitting. This prescription excludes the situation in which the slopes of the functions $\tilde{U}(\beta)|_{L_{1}}$ and $\tilde{U}(\beta)|_{L_{2}}$ are almost the same, so that the noisy data of the intersection point are avoided. Furthermore, when we choose $L_{\mathrm{min}}=8$ and $10$ for $\lambda=0$ and $8$, respectively, the resulting $\beta_{c}$ is almost independent of $L_{\mathrm{max}}$ up to the statistical error. Thus, we determine the best-fit values of $\beta_{c}$ with these choices of $L_{\mathrm{min}}$ and $L_{\mathrm{max}}=40$, where the statistical errors are under control and $\chi^{2}/\mathrm{dof}\sim 1$.
@@
Figure 5: The fitting results of the scaling behavior of the intersections $(\beta_{\times},U_{\times})$ are shown for $\lambda=0$ (left) and 88 (right). The numerical results of $\beta_{\times}$ (top) and $U_{\times}$ (bottom) are plotted against the functions $f(L_{1},L_{2})$ and $g(L_{1},L_{2})$, respectively, where the best-fit values of $y_{t}$ and $\omega$ are used to define these functions. Note that the data points are aligned on straight lines, which indicates the validity of the fitting ansatz. Here, the data symbols denote the same ones by replacing $L$ to $L_{1}$ shown in the legend of Figure 3.
@@
In Figure 5, $\beta_{\times}$ and $U_{\times}$ are plotted against $f(L_{1},L_{2})$ and $g(L_{1},L_{2})$ defined in Eq. (27), respectively, so that the linearity should indicate our finite-size scaling ansatz is working well. Here, to compute $f(L_{1},L_{2})$ and $g(L_{1},L_{2})$, we use the best-fit values of $y_{t}$ and $\omega$, obtained with $L_{\mathrm{min}}=8$ ($10$) for $\lambda=0$ (88), in Table 2. The numerical data seen in Figure 5 shows visible linearity for both $\beta_{\times}$ and $U_{\times}$ with $\lambda=0$ and $8$, suggesting the validity of our ansatz.
@@
Figure 6: Comparison of the critical temperature $\beta_{c}$ of the Heisenberg model obtained in this paper and the previous studies by the Monte Carlo [26, 27, 17, 18, 19, 20].
@@
Given the success of the finite-size scaling ansatz, we obtain the critical temperatures for the two cases of $\lambda$ from the best-fit results as
$\beta_{c}=\begin{cases}0.693035(11)\left({}^{+18}_{-20}\right)&(\lambda=0),\\ 0.319844(18)\left({}^{+43}_{-41}\right)&(\lambda=8).\end{cases}$ (31)
Here, the first parentheses denote the statistical error, which is given by the variance of the best-fit values for the 100 Jackknife samples. The second parentheses represent the systematic error from the uncertainty of the finite-size effect, which is estimated by the differences in $\beta_{c}$ for changing $L_{\mathrm{min}}\to L_{\mathrm{min}}\pm 2$, namely the values in Table 2. We use these values of $\beta_{c}$ in the following analyses. Our result of $\beta_{c}$ for $\lambda=0$ is compared with the previous Monte Carlo studies on the Heisenberg model in Figure 6.
V.4 Critical exponent: $\nu$
@@
Figure 7: The $L$ dependence of the gradient of the Binder parameter at the critical temperature $\beta_{c}$ is plotted in log-log scale for $\lambda=0$ (left) and 88 (right). The data is obtained using the interpolating function (36) as $\tilde{U}^{\prime}(\beta_{c})$. The fitting results by $a_{0}L^{1/\nu}$ are also depicted by the solid (dashed) lines inside (outside) the fitting range, which is $L\in[12,40]$ and $[18,40]$ for $\lambda=0$ and $8$, respectively.
To compare the data with the expected scaling behavior (24), we fit the data points with a function $a_{0}L^{1/\nu}$ of $L$ as shown in Figure 7. We obtain the best-fit values $(\nu,a_{0})=(0.711(1),0.0652(4))$ and $(0.603(2),0.336(7))$ with the fitting ranges $L\in[12,40]$ and $[18,40]$, for $\lambda=0$ and $8$, respectively. These ranges are taken from typical choices in the detailed analysis below.
@@
based on the scaling property in Eq. (24). Here we include the term of $L^{-\epsilon}$ as a finite-size correction, which should appear when $L$ is small, following a similar analysis in Ref. [28, 29, 12]. In principle, the correction exponent $\epsilon$ should be related to the other critical exponents such as $\omega$ (see Ref. [29]), but the precise determination of $\epsilon$ from the fits requires more precisions of data in larger $L$ asymptotics. In this paper, as did in Refs. [28, 29, 12], we fix the value of $\epsilon$ beforehand and estimate the optimal value of $\epsilon$ by the stability of the resulting fit.
@@
Figure 8: The results of $\nu$ obtained via the fitting ansatz (33) are plotted against the lower bound $L_{\mathrm{min}}$ of the fitting range for $\lambda=0$ (left) and 88 (right). In both panels, the symbols of plus-blue, cross-green, circle-red, and square-purple depict the results with $\epsilon=\infty$, $2.0$, $3.0$, and $6.0$, respectively. The horizontal lines depict the results of the constant fitting with the error bands.
@@
$\nu=0.712(1)\left({}^{+3}_{-0}\right)\left({}^{+1}_{-1}\right)_{\beta_{c}}\qquad\text{for }\lambda=0$ (34)
for the Heisenberg model. Here, the first and second parentheses denote the statistical error by the Jackknife method and the systematic error from the uncertainty of $\varepsilon$, respectively. The third error comes from the propagation of the systematic error of $\beta_{c}$.
@@
$\nu=0.601(2)\left({}^{+0}_{-2}\right)\left({}^{+5}_{-4}\right)_{\beta_{c}}\qquad\text{for }\lambda=8$ (35)
@@
These results of $\nu$ above are compared with the results of previous studies with various methods as shown in Figure 1. Here, we compare the values of $\nu$ for $\lambda=0$, $8$, and $\infty$ obtained by the Monte Carlo [12, 11, 16, 17, 18, 19, 20, 21], the conformal bootstrap [15], the $\varepsilon$-expansion [2, 7], and the functional renormalization group [8]. For reference, the conformal bootstrap predicts $\nu=0.7117(2)$ 444 Here, the error in the bootstrap bound is not one-sigma range but rigorous [15]. and the latest Monte Carlo study [12] reports $\nu=0.71164(10)$ for the Heisenberg model. Our result at $\lambda=0$ is consistent with them. On the other hand, the result at $\lambda=8$ shows a large deviation from the analytical prediction by the 1-loop $\varepsilon$-expansion. Here, we note that the difference between the 1-loop result of the $\varepsilon$-expansion and the 3-loop result at $\lambda=0$ is larger than the one between the 1-loop one at $\lambda=0$ and $\lambda=8$. It indicates that loop corrections are sizable, at least larger than the $\lambda$-dependence.
Furthermore, let us discuss the self-consistency of our results. As shown in Table 2, we obtained the estimations of $y_{t}=1/\nu$ by analysing the intersections of the Binder parameter in Section V.3, which yields $\nu=0.706(5)$ for $\lambda=0$ with $L_{\mathrm{min}}=8$ and $\nu=0.563(4)$ for $\lambda=0$ with $L_{\mathrm{min}}=10$. Compared with these values, the results of $\nu$ obtained in this section have smaller statistical errors. The smaller statistical error is mainly because we have fewer parameters to fit by considering the derivative of the Binder parameter and fixing $\beta_{c}$. The results of the two methods agree within the error for $\lambda=0$ but not well for $\lambda=8$. One of the reasons for this discrepancy could arise mainly from the choice of $L_{\mathrm{min}}$, where we set $L_{\mathrm{min}}=10$ in Section V.3 and $L_{\mathrm{min}}\geq 16$ in this section.
V.5 Critical exponents: $\eta$ and $\gamma$
@@
Figure 10: The $L$ dependence of the magnetic susceptibility at the critical temperature is plotted in log-log scale for $\lambda=0$ (left) and 88 (right). The data is obtained by the interpolating function (36) as $\tilde{\chi}_{m}(\beta_{c})|_{L}$. The fitting results by $b_{0}L^{2-\eta}$ are also depicted by the solid (dashed) lines inside (outside) the fitting range, which is $L\in[32,40]$ and $[28,40]$ for $\lambda=0$ and $8$, respectively.
@@
We fit the data points with a function $b_{0}L^{2-\eta}$ of $L$ as shown in Figure 10 to test the scaling behavior (25). Setting the fitting ranges to $L\in[32,40]$ and $[28,40]$, we obtained the best-fit values $(\eta,b_{0})=(0.031(5),0.0486(8))$ and $(0.131(8),0.140(4))$ for $\lambda=0$ and $8$, respectively. The fitting results are shown in Figure 10, where the data agree with the simple power-law scaling for sufficiently large $L$.
Figure 11: The results of $\eta$ obtained via the fitting ansatz (37) are plotted against the lower bound $L_{\mathrm{min}}$ of the fitting range for $\lambda=0$ (left) and 88 (right). In both panels, the symbols of, plus-blue, cross-green, and circle-red depict the results with $\epsilon=\infty$, $1.0$, and $2.0$, respectively. The square-purple symbols correspond to $\epsilon=4.0$ in the left panel and $6.0$ in the right panel. The colored horizontal lines depict the results of the constant fitting with the error bands.
We estimate the systematic error from the finite size effect and determine the critical exponent $\eta$ more precisely. Adding the finite size correction term to Eq. (25), we consider the scaling function to be
@@
We take the result of $\epsilon=2.0$ as the central value and use the others to estimate the systematic errors from the uncertainty of $\varepsilon$, which results in
$\eta=0.0318(27)\left({}^{+10}_{-7}\right)\left({}^{+27}_{-28}\right)_{\beta_{c}}\qquad\text{for }\lambda=0$ (38)
for the Heisenberg model. Here, the first, second, and third parentheses denote the statistical error, the systematic error from $\varepsilon$, and the systematic error from $\beta_{c}$, respectively. For reference, the conformal bootstrap predicts $\eta=0.0379(1)$ [15]. Here, the error in the bootstrap bound is not one-sigma range but rigorous.
@@
$\eta=0.132(8)\left({}^{+0}_{-4}\right)\left({}^{+12}_{-11}\right)_{\beta_{c}}\qquad\text{for }\lambda=8$ (39)
@@
$\gamma=\begin{cases}1.4009(32)\left({}^{+58}_{-0}\right)_{\nu}\left({}^{+5}_{-7}\right)_{\eta}\left({}^{+5}_{-7}\right)_{\beta_{c}}&(\lambda=0),\\ 1.123(6)\left({}^{+0}_{-3}\right)_{\nu}\left({}^{+2}_{-0}\right)_{\eta}\left({}^{+16}_{-14}\right)_{\beta_{c}}&(\lambda=8).\end{cases}$ (40)
Here, the first parentheses represent the statistical error while the second, third, and fourth denote the systematic errors from $\nu$, $\eta$, and $\beta_{c}$, respectively.
Figure 12: The critical exponent $\gamma$ obtained in this paper and the previous studies by the Monte Carlo [12, 11, 16, 17, 18, 19, 20] (colored open symbols), the conformal bootstrap [15] (star-yellow), the $\varepsilon$-expansion [7] (filled symbols) are compared. The filled circle-black symbols depict the 1-loop results of the $\varepsilon$-expansion in Ref. [2] and the filled square-circle ones are the 3-loop results in Ref. [7].
Our results of $\gamma$ are compared with the results of the previous Monte Carlo studies [12, 11, 16, 17, 18, 19, 20], the $\varepsilon$-expansion [2, 7], and conformal bootstrap [15] in Figure 12. The 3-loop calculations yield coincidental results $\gamma=1.380(8)$ and $1.381(8)$ for the Heisenberg model ($\lambda=0$) and the Heisenberg-dipolar model with $\lambda=\infty$ [31, 5]. Also, the conformal bootstrap predicts $\gamma=1.3964(5)$ [15]. On the other hand, our numerical results of $\gamma$ show a clear difference between $\lambda=0$ and $8$. This observation suggests a possibility that the difference between the critical exponents at the dipolar and Heisenberg fixed points is more significant than the expectation from the 3-loop calculation. Although our numerical method has room to be improved for more rigorous investigation, our result is consistent with the previous studies at least for $\lambda=0$. For instance, the latest one, $\gamma=1.39635(20)$, obtained by the Monte Carlo study [12] is $1.4\sigma$ consistent(see the detailed comparison in Figure 12).
VI Correlation functions and transverse suppression